OpenStudy (zolock):

Find the constant of variation for the quadratic variation. (please help fast I need to turn it in) x=2,3,4,5,6 y=24,54,96,150,216 a=12 b=6 c=30 d=18

OpenStudy (whpalmer4):

if you have quadratic variation, then \(y = kx^2\) where \(k\) is the constant of variation. \[24=k(2)^2\]\[k=\]\[54=k(3)^2\]\[k=\]\[96=k(4)^2\]\[k=\]etc.